There are 10 defective light bulbs in a bin containing 20 light bulbs. In how many ways can you select at most 2 defective light bulbs when you choose 10 light bulbs at random?


1
Expert's answer
2022-03-04T09:25:09-0500

Let "X=" the number of defective light bulbs in the sample.


"X=0: \dbinom{10}{0}\dbinom{10}{10-0}=1(1)=1"

"X=1: \dbinom{10}{1}\dbinom{10}{10-1}=10(10)=100"

"X=2: \dbinom{10}{2}\dbinom{10}{10-2}=45(45)=2025"

"1+100+2025=2126"

We can select at most 2 defective light bulbs in 2126 ways.